# Simultaneous Task Allocation and Planning for Multi-Robots under Hierarchical Temporal Logic Specifications

### v3

@@ VIII-B2 Scenario 2
Distribute documents to desks $d_{10}$, $d_{7}$, and $d_{5}$, and avoid public areas while carrying the document.
@@ VIII-B3 Scenario 3
Take a photo in meeting rooms $m_{1}$, $m_{4}$, and $m_{6}$. The camera should be turned off for privacy reasons when not in meeting rooms. Deliver a document from desk $d_{5}$ to $d_{3}$, ensuring it does not pass through any public areas, as the document is internal and confidential. Guide a person waiting at desk $d_{11}$ to meeting room $m_{6}$.
@@ VIII-B4 Combinations of scenarios 1, 2 and 3
We examine combinations of any two of these tasks as well as the combination of all three. Due to space constraints, we only detail the scenario involving the combination of all three tasks below.
@@
$L_{2}:\quad$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{1}=\Diamond\phi_{3}^{1}\wedge\Diamond\phi_{3}^{2}}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{2}=\Diamond\phi_{3}^{3}\wedge\Diamond\phi_{3}^{4}\wedge\Diamond\phi_{3}^{5}}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{3}=\Diamond\phi_{3}^{6}\wedge\Diamond\phi_{3}^{7}\wedge\Diamond\phi_{3}^{8}}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{2}=\Diamond(d_{5}\wedge\mathsf{emptybin}\wedge\bigcirc(d_{5}\wedge\mathsf{default})})$
@@
scenario $l_{\text{flat}}$ $l_{\text{hier}}$ ${\mathcal{A}}_{\text{flat}}$ ${\mathcal{A}}_{\text{hier}}$ $t_{\text{flat}}$ $t_{\text{hier}}$ $c_{\text{flat}}$ $c_{\text{hier}}$
@@
1 $\wedge$ 2 $\wedge$ 3 79 111 timeout (70, 112) timeout 89.6$\pm$28.1 246.6$\pm$8.3
TABLE III: The comparative analysis focuses on two different types of sc-LTL specifications. We denote the lengths of the flat and hierarchical sc-LTL specifications as $l_{\text{flat}}$ and $l_{\text{hier}}$, respectively. The sizes of the corresponding NBAs are represented by ${\mathcal{A}}_{\text{flat}}$ and ${\mathcal{A}}_{\text{hier}}$, which detail the number of nodes and edges, with the node count listed first. In terms of solutions, the runtimes for the flat and hierarchical sc-LTL specifications are indicated by $t_{\text{flat}}$ and $t_{\text{hier}}$, respectively. Additionally, the plan horizons, or the lengths of the solutions, for the flat and hierarchical specifications are denoted by $c_{\text{flat}}$ and $c_{\text{hier}}$, respectively.
@@ VIII-D Comparison with Existing Works
We use $m=6$ robots and compare our method with the approach in [8], both using all heuristics. Note that [8]’s method is tailored to flat sc-LTL specifications. Due to the absence of open-source code, we implemented their method. Robot locations in each scenario are randomly assigned within the free space. The performance, in terms of average runtimes and costs over 20 runs, is detailed in Tab. III and includes the length of formulas and sizes of automatons. The length of a formula is the total number of logical and temporal operators. We observed that, contrary to simulation findings in [42], where hierarchical sc-LTL shows syntactical conciseness compared to flat sc-LTL, our hierarchical sc-LTL specifications are longer in length. We hypothesize that the syntactical succinctness of hierarchical sc-LTL becomes more evident in scenarios with complex interdependencies between tasks at non-bottom levels, which are more challenging to express through flat sc-LTL specifications. This is not the case in our scenarios, where tasks at non-bottom levels are temporally independent. Upon reviewing the results, [8]’s method failed to generate solutions for the last four tasks within the one-hour limit. For tasks 1 and 2, our method produced solutions more quickly and with comparable costs. The failure of  [8]’s approach is attributed to the excessively large automatons it generates, sometimes with hundreds of thousands of edges, e.g., 325745 edges for scenario $2\wedge 3$, making the computation of the decomposition set time-consuming as it requires iterating over all possible runs. For the most complex scenario $1\wedge 2\wedge 3$, generating an automaton within one hour is impossible. In contrast, our method was able to find a solution in around 90 seconds.
@@ VIII-E Effects of Granularity of Hierarchy
For the same task, the performance of the planning algorithm is influenced by the granularity of hierarchical specifications, given the variability in humans’ levels of hierarchical reasoning granularity. To investigate this, we evaluate the most complex task, which combines scenarios 1, 2, and 3 as defined in (VIII-B4). The specifications in (VIII-B4) represent the finest granularity. We then progressively increase the coarseness by flattening more non-leaf specifications through the merging of their respective leaf specifications, thus gradually transforming the hierarchical structure towards a more flattened form. Eventually, this process results in a single flat specification. For example, flattening specification $\phi_{3}^{6}$ results in $\phi_{3}^{6}=\Diamond(m_{1}\wedge\mathsf{photo})\wedge\Diamond(m_{4}\wedge\mathsf{photo})\wedge\Diamond(m_{6}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})$. The outcomes from 20 runs are presented in Tab. IV. Generally, as more non-leaf specifications are flattened, the automatons increase in size, resulting in longer runtimes, while the costs remain relatively stable. Consequently, finer granularity in the hierarchy is associated with enhanced planning performance.
@@ VIII-F Scalability
Figure 10: The presented keyframes illustrate the execution of the derivative task described in Eq. (VIII-G) within the AI2-THOR framework, indicating the crucial phases in the completion of task. Each of the four robots is inside yellow boxes with distinct numbers. The upper-left corner of the keyframes marks the time steps, and the objects being manipulated are encircled in blue. Note that the robots 2 and 4 are simultaneously manipulating the box and the laptop.
@@ VIII-G Evaluation on simulator AI2-THOR
We conduct our experiments using the AI2-THOR simulator and the ALFRED dataset. AI2-THOR [70] provides a realistic 3D simulated environment for robots to perform interactions within various household settings such as kitchens and living rooms. This setting allows for the assessment of algorithms in tasks that include navigation and object manipulation. The ALFRED dataset [71], used within AI2-THOR, focuses on natural language instructions for simple, sequentially ordered tasks. Examples of such tasks in a living room setting include (i) “Turn off the floor lamp,” (ii) “Move the box to a sofa,” (iii) “Move the laptop from the table to an armchair,” and (iv) “Dispose of the credit card in the garbage can.” We refer to these tasks as base tasks. Then we construct derivative tasks by combining several base tasks within the same floor plan. One derivative task from base tasks (i)-(iv) is that “First turn off the floor lamp. After that, move the box to a sofa and the laptop to an armchair. Finally, dispose of the credit card in the garbage can.” The hierarchical sc-LTL formulas are:
@@
$J(\tau)=\omega\max_{r\in[N]}c_{r}+(1-\omega)\sum_{r=1}^{N}c_{r}$ (11)
where $\omega\in[0,1]$. Here, setting $\omega=0$ simplifies (11) to objective (3), while $\omega=1$ seeks to reduce the maximum cost among robots, with the secondary term regularizing the costs of the remaining robots. We set $\omega=0.9$ to prioritize minimizing the maximum path length among the robots. The results, detailed in Table V, indicate an increase in runtime for solutions as the number of base tasks or robots increases. Additionally, with more robots engaged, the maximum cost per robot is reduced due to the ability to distribute tasks more evenly among multiple robots. The execution of derivative task (VIII-G) is illustrated in Fig. 10.
# base tasks # robots $t$ $c$
1 1 2.6$\pm$0.9 11.7$\pm$8.7
2 3.0$\pm$1.2 11.7$\pm$9.8
4 4.1$\pm$1.3 9.0$\pm$7.6
2 1 4.3$\pm$2.9 23.4$\pm$21.0
2 4.8$\pm$3.2 17.1$\pm$15.4
4 6.5$\pm$4.7 14.8$\pm$12.7
3 1 10.8$\pm$7.7 38.2$\pm$25.4
2 12.1$\pm$9.2 27.9$\pm$21.6
4 14.6$\pm$9.1 22.0$\pm$14.2
4 1 13.0$\pm$13.9 50.1$\pm$29.6
2 13.3$\pm$14.9 30.7$\pm$16.9
4 14.9$\pm$11.4 23.9$\pm$14.0
TABLE V: Results comparing performance across different numbers of base tasks and robots. Here, $t$ represents the runtimes required to return solutions, and $c$ denotes the maximal cost incurred by any robot.
@@ IX Conclusion and Discussion
In this work, we addressed the problem of STAP for multiple robots subject to hierarchical sc-LTL specifications. First, we presented the syntax and semantics for hierarchical sc-LTL. Next, we developed a search-based planning algorithm. To further expedite the search, we introduced several heuristics leveraging the task structure. We also provided theoretical analysis on the completeness and optimality.
Our simulations, focused on service tasks involving navigation and manipulation, included an ablation study on heuristics. This study demonstrated that each heuristic variably accelerates the search, with their combination yielding a significant speedup. Comparative studies with existing works illustrate that our approach can handle complex tasks beyond the reach of current methods. Scalability tests reveal that our method can manage up to 30 robots within 200 seconds. We attribute the computational efficiency to two factors. First, the adoption of hierarchical specifications results in smaller automatons, as indicated in Tab. III. We contend that hierarchical specifications do not necessarily increase cognitive load, given that humans naturally excel at hierarchical thinking [6, 7]. The results from the user study also highlighted the user-friendliness of hierarchical sc-LTL. The second factor is the design of the planning algorithm, particularly its heuristics. As shown in Tab. II, the planning algorithm performs faster with any heuristic than without. Furthermore, employing a combination of all three heuristics results in shorter runtimes compared to using any single heuristic.
While the hierarchical sc-LTL introduced has proven effective, there are still several areas that require additional investigation. Firstly, the planning algorithm assumes that robots operate independently, without considering scenarios where coordination, such as jointly carrying heavy items, is necessary. Therefore, it is crucial to include such situations in the algorithm. Another area of interest is addressing uncertainty arising from multiple sources, including partially observable environments, or the costs of actions, which might not be predetermined. Since hierarchical sc-LTL can encompass multiple conflicting specifications, a future research avenue could be to define and maximize degree of satisfaction when all specifications cannot be concurrently met. Lastly, integrating natural languages, as in work [72], to embed task hierarchies from human instructions into structured planning languages suitable for existing planners, presents an intriguing prospect.

### v4

@@ VIII-B2 Scenario 2
Distribute documents to desks $d_{10}$, $d_{7}$, and $d_{5}$, and avoid public areas while carrying the document. Let $\mathsf{carry}$ denote the action of the robot carrying the document.
@@ VIII-B3 Scenario 3
Take a photo in meeting rooms $m_{1}$, $m_{4}$, and $m_{6}$. The camera should be turned off for privacy reasons when not in meeting rooms. Deliver a document from desk $d_{5}$ to $d_{3}$, ensuring it does not pass through any public areas, as the document is internal and confidential. Guide a person waiting at desk $d_{11}$ to meeting room $m_{6}$. Let $\mathsf{guide}$, $\mathsf{photo}$, and $\mathsf{camera}$ denote the actions of the robot guiding a person, capturing a photo, and activating the camera, respectively.
@@ VIII-B4 Combinations of scenarios 1, 2 and 3
We examine combinations of any two of these tasks as well as the combination of all three. Due to space constraints, we only detail the scenario involving the final occurrence of all three tasks.
@@
[…]
@@
$\phi_{3}^{2}=\Diamond(g\wedge\bigcirc(g\wedge\mathsf{emptybin})\wedge\Diamond(d_{5}\wedge\bigcirc(d_{5}\wedge\mathsf{default})))$
@@
If the objective is to accomplish either scenario, this can be indicated by replacing the formula at level $L_{1}$ with $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\vee\phi_{2}^{2}\vee\phi_{2}^{3})$. Furthermore, the expression $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\vee\phi_{2}^{2})\wedge\Diamond\phi_{2}^{3}$ specifies that either task 1 or task 2 must be completed, in addition to task 3.
scenario $l_{\text{std}}$ $l_{\text{hier}}$ ${\mathcal{A}}_{\text{std}}$ ${\mathcal{A}}_{\text{hier}}$ $t_{\text{std}}$ $t_{\text{hier}}$ $c_{\text{std}}$ $c_{\text{hier}}$
@@
1 $\wedge$ 2 $\wedge$ 3 79 111 (70, 112) timeout 89.6$\pm$28.1 246.6$\pm$8.3
(1 $\vee$ 2) $\wedge$ 3 79 110 (66, 98) timeout 71.5$\pm$17.6 213.1$\pm$23.4
1 $\vee$ 2 $\vee$ 3 79 109 (64, 94) timeout 63.4$\pm$20.9 163.9$\pm$40.5
TABLE III: The comparative analysis focuses on two different types of LTLf specifications. We denote the lengths of the standard and H-LTLf specifications as $l_{\text{std}}$ and $l_{\text{hier}}$, respectively. The sizes of the corresponding NFAs are represented by ${\mathcal{A}}_{\text{std}}$ and ${\mathcal{A}}_{\text{hier}}$, which detail the number of nodes and edges, with the node count listed first. In terms of solutions, the runtimes for the standard and H-LTLf specifications are indicated by $t_{\text{std}}$ and $t_{\text{hier}}$, respectively. Additionally, the plan horizons, or the lengths of the solutions, for the standard and hierarchical specifications are denoted by $c_{\text{std}}$ and $c_{\text{hier}}$, respectively.
@@ VIII-D Comparison with Existing Works
We use $m=6$ robots and compare with the approach in [9], both using all heuristics. Note that [9]’s method is tailored to standard LTLf specifications. Due to the absence of open-source code, we implemented their method. Robot locations in each scenario are randomly assigned within the free space. The performance, in terms of average runtimes and costs over 20 runs, is detailed in Tab. III and includes the length of formulas and sizes of automata. The length of a formula is the total number of logical and temporal operators. Upon reviewing the results, [9]’s method failed to generate solutions for the last seven tasks within the one-hour limit. For tasks 1 and 2, our method produced solutions more quickly and with comparable costs. The failure of  [9]’s approach is attributed to the excessively large automata it generates, sometimes with hundreds of thousands of edges, e.g., 325745 edges for scenario $2\wedge 3$, making the computation of the decomposition set time-consuming as it requires iterating over all possible runs. For the most complex scenario, such as $1\wedge 2\wedge 3$, generating an automaton within one hour is impossible. In contrast, our method was able to find a solution in around 90 seconds. Moreover, considering the last three tasks, which require completing all, two, or only one task, both the runtime and cost decrease as the number of required tasks is reduced.
@@ VIII-E Effects of Granularity of Hierarchy
For the same task, the performance of the planning algorithm is influenced by the granularity of hierarchical specifications. To investigate this, we evaluate the most complex task, which combines scenarios 1, 2, and 3 as defined in (VIII-B4). The specifications in (VIII-B4) represent the finest granularity. We then progressively increase the coarseness by flattening more non-leaf specifications through the merging of their respective leaf specifications, thus gradually transforming the hierarchical structure towards a more flattened form. Eventually, this process results in a single standard specification. For example, flattening specification $\phi_{3}^{6}$ results in $\phi_{3}^{6}=\Diamond(m_{1}\wedge\mathsf{photo})\wedge\Diamond(m_{4}\wedge\mathsf{photo})\wedge\Diamond(m_{6}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})$. The outcomes from 20 runs are presented in Tab. IV. Generally, as more non-leaf specifications are flattened, the automata increase in size, resulting in longer runtimes, while the costs remain relatively stable. Consequently, finer granularity in the hierarchy is associated with enhanced planning performance.
@@ VIII-F Scalability
Figure 11: The keyframes illustrate the execution of the derivative task described in Eq. (VIII-G) within the AI2-THOR framework, indicating the crucial phases in the completion of task. Each of the four robots is inside yellow boxes with distinct numbers. The upper-left corner of the keyframes indicates the time instants, while the bottom-left corner displays the working robot and its corresponding action. The working robots are enclosed in blue boxes, with dashed lines in a magenta gradient color connecting the position where they performed most recent manipulation action to their current location. White arrows highlight the objects with which the robots are interacting. Robots 2 and 4 are simultaneously manipulating the box and the laptop.
@@ VIII-G Evaluation on simulator AI2-THOR
We conduct our experiments using the AI2-THOR simulator and the ALFRED dataset. AI2-THOR [69] provides a realistic 3D simulated environment for robots to perform interactions within various household settings such as kitchens and living rooms. This setting allows for the assessment of algorithms in tasks that include navigation and object manipulation. The ALFRED dataset [70], used within AI2-THOR, focuses on natural language instructions for simple, sequentially ordered tasks. Examples of such tasks in a living room setting include (i) “Turn off the floor lamp,” (ii) “Move the box to a sofa,” (iii) “Move the laptop from the table to an armchair,” and (iv) “Dispose of the credit card in the garbage can.” We refer to these tasks as base tasks. Then we construct derivative tasks by combining several base tasks within the same floor plan. One derivative task from base tasks (i)-(iv) is that “First turn off the floor lamp. After that, move the box to a sofa and the laptop to an armchair. Finally, dispose of the credit card in the garbage can.” The H-LTLf formulas are:
@@
$J(\tau)=\omega\max_{r\in[N]_{+}}c_{r}+(1-\omega)\sum_{r=1}^{N}c_{r}$ (12)
where $\omega\in[0,1]$. Here, setting $\omega=0$ simplifies (12) to objective (4), while $\omega=1$ seeks to reduce the maximum cost among robots, with the secondary term regularizing the costs of the remaining robots. In the evaluation, we set $\omega=0.9$ to prioritize minimizing the completion time (makespan). The results, detailed in Table V, indicate an increase in runtime for solutions as the number of base tasks or robots increases. Additionally, with more robots engaged, the completion time is reduced due to the ability to distribute tasks more evenly among multiple robots. The execution of derivative task (VIII-G) is illustrated in Fig. 11.
# robots # tasks $t$ $c$ tasks $t$ $c$
1 1 2.6$\pm$0.9 11.7$\pm$8.7 3 10.8$\pm$7.7 38.2$\pm$25.4
2 3.0$\pm$1.2 11.7$\pm$9.8 12.1$\pm$9.2 27.9$\pm$21.6
4 4.1$\pm$1.3 9.0$\pm$7.6 14.6$\pm$9.1 22.0$\pm$14.2
1 2 4.3$\pm$2.9 23.4$\pm$21.0 4 13.0$\pm$13.9 50.1$\pm$29.6
2 4.8$\pm$3.2 17.1$\pm$15.4 13.3$\pm$14.9 30.7$\pm$16.9
4 6.5$\pm$4.7 14.8$\pm$12.7 14.9$\pm$11.4 23.9$\pm$14.0
TABLE V: Results across different numbers of base tasks and robots. Here, $t$ represents the runtimes required to return solutions, and $c$ denotes the completion time.
@@ IX Conclusion and Discussion
We addressed the problem of STAP for multiple robots subject to H-LTLf specifications. First, we presented the syntax and semantics for H-LTLf. Next, we developed a search-based planning algorithm, and introduced several heuristics leveraging the task structure to expedite it. We provided theoretical analysis on the completeness and optimality.
Our simulations, focused on service tasks involving navigation and manipulation, included an ablation study on heuristics. This study demonstrated that each heuristic variably accelerates the search, with their combination yielding a significant speedup. Comparative studies with existing works illustrate that our approach can handle complex tasks beyond the reach of current methods. Scalability tests reveal that our method can manage up to 30 robots within 200 seconds. We attribute the computational efficiency to two factors. First, the adoption of hierarchical specifications results in smaller automatons, as indicated in Tab. III. The results from the user study also highlighted the user-friendliness of H-LTLf. The second factor is the planning algorithm, particularly its heuristics. As shown in Tab. II, the planning algorithm performs faster with any heuristic than without. Furthermore, employing a combination of all three heuristics results in shorter runtimes compared to using any single heuristic.
While the H-LTLf introduced has proven effective, there are several areas that require investigation. An open challenge is developing an algorithmic approach to convert standard LTLf formulas into H-LTLf formulas and exploring how to introduce the hierarchical form to full LTL. The second aspect relates to the planning algorithm. Currently, the planning algorithm assumes that robots operate independently, without considering scenarios where coordination, such as jointly carrying heavy items, is necessary. Therefore, it is crucial to include such situations, as in work [71]. Since H-LTLf can encompass multiple conflicting specifications, a future direction could be to define and maximize degree of satisfaction when not all specifications can be met. Finally, integrating natural languages, as in work [72], to embed task hierarchies from instructions into structured formal languages, presents an intriguing prospect.